OpenStudy (anonymous):

Ky has 3 times more books than Grant, and Grant has 6 fewer books than Jaime. if the total combined number of books is 176, how many books does Jaime have? A) 28 B) 34 C) 40 D) 102

OpenStudy (anonymous):

2 + 2 = 5

OpenStudy (anonymous):

I attempted it and I didn't get the forming of the equation correct but here is what I did 3x(-6+x)=176

OpenStudy (anonymous):

um

OpenStudy (anonymous):

you sure this is algebra

OpenStudy (anonymous):

yup, it's just placed in confusing word problems :/

OpenStudy (lastdaywork):

Given K = 3G J - 6 = G K + G + J = 176 Can you find the value of J ??

OpenStudy (anonymous):

so substitute so it becomes, 3(j-6)+j-6+j=176 ??

OpenStudy (lastdaywork):

Just a small problem; J should be a natural number...but the answer is coming out to be a fraction :(

hartnn (hartnn):

j = 200/5 = 40 its not a fraction

OpenStudy (lastdaywork):

Okay..I made a calculation mistake..my bad XD

OpenStudy (anonymous):

do I get a medal?